# More than 200 Globular Clusters in the Milky Way, and still no super-Solar metallicity ones

### v1

@@ 7 Conclusion
Several dozens of new Galactic GCs have been discovered in the past dozen years thanks to the advent of new large surveys in the optical and near-IR, significantly augmenting the MW GC system. Therefore, we wanted to collect them in a single catalogue, in order to show that in reality our Galaxy could host more than 200 GCs. Taking into account all the limitations expressed in Section 2, this work allowed us to upgrade the MW GC system, adding 37 new GC candidates, also we were able to upgrade the LF and MD. One of the main achievements of this work is the fact that we are exploring the faint tail of the MW GC LF. The ”upgraded” GC LF is shifted towards slightly lower luminosities, and it may indicate a double-peaked distribution. When comparing the MW GC LF with that of Andromeda’s GC LF, and assuming the authenticity of the second peak, it suggests that the MW may harbour accreted objects. Alternatively, these objects could be survivors of a dynamical process that led to their dissolution or disruption. Due to the fact that their age estimates are approximate, we can only speculate on the fact that some of these low-luminosity objects could be what remains of those objects that today the James Webb Space Telescope is observing at high redshift (see Section 1), or there may be a connection between those compact red objects and diffuse Galactic GCs, present in our compilation. Moreover, the MD shows a bimodal distribution, with a third tentative peak. We find that new GCs follow the same behaviour as the bulge GC population (as shown in Figure 6), as expected due to their location inside the MW. Furthermore, it seems to be accepted that bulge GCs may be the first stellar objects formed during the formation of the bulge and/or disk (an example may be Liller 1 - Dalessandro et al. 2022), while the metal-poor GCs, located at higher Galactocentric distances, may represent the accreted component (e.g. Forbes & Bridges 2010 and reference therein). Nevertheless, we cannot exclude that the Galactic bulge (as well as the Galactic halo) may contain ex-situ clusters or those formed in a second moment from accreted gas, supporting the scenario that the bulge was formed first by monolithic collapse, and then grown up hierarchically by merging galaxies. This is a hypothesis that must be confirmed spectroscopically. Furthermore, no GCs with super-solar metallicity have been discovered so far. This is perhaps the most important conclusion. In order to give clues about the so many questions that we wondered in Section 5, we definitely need the support of (i) the spectroscopic analysis in order to confirm the nature of these GCs, derive their chemistry and reconstruct their orbits; (ii) numerical simulations that can help to quantify how many super metal-rich GCs we can expect in the MW, if they can or cannot exist, if anything they were destroyed, understand why; and finally (iii) deeper observations in order to reach the MSTO and derive robust ages.
@@
This is the outcome of existing surveys, such as the VVV-VVVX, 2MASS, SDSS in the near-infrared (NIR), and Gaia mission in the optical wavelengths. Hence, more progress is expected to come with the advent of future instrumentations (LSST, MOONS, 4MOST, just to mention a few).

### v2

@@ 8 Conclusion
Several dozens of new Galactic GCs have been discovered in the past decade or so thanks to the advent of new large surveys in the optical and near-IR, significantly augmenting the MW GC system. Therefore, our aim is to collect them in a single catalog to show that, in reality, our Galaxy could potentially host more than 200 GCs. Taking into account all the limitations described in Section 2, this work allowed us to update the data on the MW GC system, adding 37 new GC candidates; in addition, we were able to upgrade the LF and MD data. Independently, Bica et al. (2024) analyzed a similar updated sample of GCs in the bulge, arriving at analogous conclusions.
@@
One of the main achievements of our work is the fact that we are exploring the faint tail of the MW GC LF. The ”updated” GC LF is shifted toward slightly lower luminosities and may indicate a double-peaked distribution. When comparing the MW GC LF with that of Andromeda’s GC LF and assuming the authenticity of the second peak, it suggests that the MW may harbour accreted objects. Alternatively, these objects could be survivors of a dynamical process that led to their dissolution or disruption. Due to the fact that their age estimates are approximate, we can only speculate on the fact that some of these low-luminosity objects could be what remains of those objects that today the James Webb Space Telescope is observing at high redshift (see Section 1). Alternatively, there may be a connection between those compact red objects and diffuse Galactic GCs, present in our compilation. Moreover, the MD shows a bimodal distribution, with a third tentative peak. We find that new GCs follow the same behavior as the bulge GC population (as shown in Figure 6), as expected due to their location inside the MW. Furthermore, it seems to be accepted that bulge GCs may be the first stellar objects formed during the formation of the bulge and/or disk (an example may be Liller 1 - Dalessandro et al. 2022), while the metal-poor GCs, located at higher Galactocentric distances, may represent the accreted component (e.g., Forbes & Bridges 2010 and reference therein). Nevertheless, we cannot exclude the possibility that the Galactic bulge (as well as the Galactic halo) may contain ex situ clusters or those formed in a second moment from accreted gas, supporting the scenario that the bulge was formed first by monolithic collapse and then grew up hierarchically via merging galaxies. This is a hypothesis that must be confirmed spectroscopically. Furthermore, no GCs with super-solar metallicity have been discovered so far. This is perhaps the most important conclusion. To obtain clues to resolving the questions we consider in Section 5, we definitely need the support of: (i) spectroscopic analyses to confirm the nature of these GCs, derive their chemistry, and reconstruct their orbits; (ii) numerical simulations that can help quantify how many super metal-rich GCs we can expect in the MW, if they can or cannot exist, if they were destroyed, and if so, to understand why and how; and, finally, (iii) deeper observations in order to reach the MSTO and derive robust ages.
@@
The data and conclusions presented in this work are compiled on the basis of existing surveys, such as VVV-VVVX, 2MASS, and SDSS in the near-infrared (NIR), as well as the Gaia mission in the optical wavelengths. Hence, more progress is expected to come with the advent of future instrumentations such as LSST, MOONS, 4MOST, and others.